OpenStudy (anonymous):

If x squared + y squared=36 and (x+2) quantity squared=64, what is the value of xy?

OpenStudy (anonymous):

ahhh

hartnn (hartnn):

can u find x from \((x+2)^2=64\) ??

OpenStudy (anonymous):

14

OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

for question 1.

OpenStudy (anonymous):

cause

OpenStudy (anonymous):

x squared plus y squared=36

OpenStudy (anonymous):

2xy has to equal 28

OpenStudy (anonymous):

so half of 28, which is half of 2xy,=14

OpenStudy (anonymous):

so the answer is B.

hartnn (hartnn):

how 2xy = 28 ? u had x^2+y^2 ?

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

x squared plus y squared=36

OpenStudy (anonymous):

(x+y) quantity squared=64

OpenStudy (anonymous):

it includes the number 2xy

OpenStudy (anonymous):

which has to equal 28

OpenStudy (anonymous):

half ot 2xy=14

OpenStudy (anonymous):

xy=14

hartnn (hartnn):

is 0 in the option ?

OpenStudy (anonymous):

36+28=64

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

the answer booklet says B.

OpenStudy (anonymous):

x times y =14

OpenStudy (anonymous):

do you understand where I am getting at?

hartnn (hartnn):

hmm....maybe question is incorrect....

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):

look it

OpenStudy (anonymous):

x squared plus 2xy plus y squared=64

OpenStudy (anonymous):

x squared plus y squared=36

OpenStudy (anonymous):

36+2xy=64

OpenStudy (anonymous):

2xy=28

OpenStudy (anonymous):

xy=14

OpenStudy (anonymous):

the value of xy=14

OpenStudy (anonymous):

therefore, the answer is B.

OpenStudy (anonymous):

do you understand where I am getting at?

OpenStudy (anonymous):

Where are you getting 2xy? You only get 2xy if the question is: (x + y)^2, not x^2 + y^2 (like your question says).

OpenStudy (anonymous):

look it

OpenStudy (anonymous):

the does say (x+y) quantity squared=64

OpenStudy (anonymous):

the ? does

OpenStudy (anonymous):

x squared+y squared=36

OpenStudy (anonymous):

(x+y) quantity squared=64

OpenStudy (anonymous):

do you see now?

hartnn (hartnn):

u wrote x+2 ....

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

it meant to write (x+y) quantity squared

OpenStudy (anonymous):

so

OpenStudy (anonymous):

as I was saying

OpenStudy (anonymous):

36+2xy=64

OpenStudy (anonymous):

2xy+28

OpenStudy (anonymous):

xy=14

hartnn (hartnn):

then u are correct....

OpenStudy (anonymous):

xy=14

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

next problem :)